AN AB INITIO MOLECULAR ORBITAL STUDY OF THE GROUND AND LOW-LYING ELECTRONIC EXCITED STATES OF FeN: PREDICTION TOWARD MOLECULAR SPECTROSCOPY

SACHIKO S. ITONO and \underline{TSUNEO HIRANO}, Department of Chemistry, Faculty of Science, Ochanomizu University, Tokyo 112-8610, Japan; UMPEI NAGASHIMA, Tsukuba Advanced Computing Center, National Institute of Advanced Industrial Science and Technology, Ibaraki 305-8562, Japan.

To answer the long-standing debate which state is the electronic ground state of FeN, highly correlated ab initio molecular orbital calculations with or without relativistic corrections and spin-orbit coupling interaction have been carried out for several low-lying electronic states including the ^{2}Δ, ^{4}Π, ^{4}Φ, ^{2}Π, ^{2}Φ, and ^{6}Σ^{+} states. The MR-ACPF calculations as well as the MR-SDCI + Q calculations showed that the electronic ground state of FeN is ^{2}Δ. Coinsidence between the predicted spectroscopic constants and those observed by Aiuchi et al.^{a} and Ziurys et al.^{b} shows that the state they observed is the X^{2}Δ Ω =2/5 spin substate. The equilibrium bond length r_{e}, rotational constant B_{0}, centrifugal distortion constant D_{J}, harmonic frequency ω_{e}, anharmonicity constant ω_{e}x_{e} , and spin-orbit coupling constant A predicted for the X^{2}Δ state at the MR-ACPF level are 1.5752 Å, 0.60334 cm^{−1}, 1.17 x 10^{−6} cm^{−1}, 890 cm^{−1}, 8.2 cm^{−1}, and -362 cm^{−1}, respectively.

aK. Aiuchi and K. Shibuya, J. Mol. Spectrosc. \underline{204}, 235, (2000).

bP. M. Sheridan, M. A. Brewster, and L. M. Ziurys, Proc. of 56th Ohio State Univ. Int. Symp. Molec. Spectrosc. \underline{FB06}, 207,